uxx= c^2ut by variable separation method
1
Expert's answer
2021-10-13T14:22:00-0400

We will solve the problem in next full description

1) Differential equation

"u_t=c^2u_{xx}" "0<x<l.\space t>0"

2) Boundary conditions (zero conditions for example)

u(0,t)=0, u(l,t)=0;

3) Initial condition

u(x,0)="\phi(x),0<x<l"

Solution:

Let u(x,y)=X(x)T(t) or have separeted variables.

Inserting this form in differential equation we have:

"X(x)T'(t)=c^2T(t)X''(x)"

Dividing both paths by "c^2X(x)T(t)" we have

"\frac{1}{c^2}\cdot \frac{T'(t)}{T(t)}=\frac{X''(x)}{X(x)}"

In this equation left part depend on t bur right part depend on x therefore both part are equal some constant which may be denoted as -"\lambda"

Thus we have the system:

"x = \begin{cases}
 T'+\lambda c^2T(t)=0\\
 X''(x)+\lambda\cdot X(x)=0
\end{cases}"

Let we consider second eqution

"X''(x)+\lambda\cdot X(x)=0"

Boundary conditions at x=0 and x=l take the form:

X(0)=X(l)=0.

1) "\lambda<0"

General solution of ode is "X(x)=C_1e^{-\sqrt{-\lambda}x}+C_2e^{\sqrt{\lambda }x}"

Then "X(0)=C_1+C_2=0" ,"C_2=-C_1"

Therefore "X(x)=C\cdot (e^{-\sqrt{-\lambda}x}-e^{\sqrt{-\lambda}x})"

Condition X(l)=0 implies

"C\cdot (e^{-\sqrt{-\lambda}l}-e^{\sqrt{-\lambda}l})=0" and hence C=0, therefore X(x)"\equiv0"

but we need nontrival solution hence case "\lambda<0" is impossible, case "\lambda=0" is impossible too with similar argues, typical case is "\lambda>0"

In this case "X(x)=C_1sin(\sqrt \lambda x)+C_2cos(\sqrt \lambda x)"

X(0)= "C_2=0=>" X(x)="Csin(\sqrt \lambda x)"

X(l)=0=>"sin(\sqrt\lambda l)=0=>\sqrt \lambda l=\pi\cdot n.n>0"

"\lambda_n=\left(\frac{\pi\cdot n}{l} \right)^2" and "X_n(x)=sin\left(\frac{\pi\cdot n\cdot x}{l}\right)"

For this "\lambda_n" an equation "T'+\lambda c^2T(t)=0" has solution

"T_n(t)=e^{-\lambda_n\cdot c^2\cdot t}" and "u+n(x,t)=T_n(t)X_n(x)=e^{-\lambda_nc^2t}sin\left( \frac{\pi nx}{l}\right)"

After that we search solution in the form

"\sum_{n=1}^{\infty}C_ne^{-\lambda_nc^2t}sin\left( \frac{\pi nx}{l}\right)" for initial condition to satisfy.

Insertinf t=0 we have

"\sum_{n=1}^{\infty}C_nsin\left( \frac{\pi nx}{l}\right)=\phi(x),0<x<l"

Then "C_n=\frac{2}{l}\int_0^l\phi(x)sin\frac{\pi n z}{l}dz" -Fourier coefficients. With such coefficients solution of the problem is:

"\sum_{n=1}^{\infty}C_ne^{-\lambda_nc^2t}sin\left( \frac{\pi nx}{l}\right)"